OpenStudy (anonymous):

4x + 2y = 12 x – y = 3

OpenStudy (anonymous):

No solutions.

OpenStudy (anonymous):

Solve the second equation for x: x = 3 + y Substitute this x in for the x in the first equation: 4(3 + y) + 2y = 12 12 + 4y + 2y = 12 6y = 0 y = 0 Plug this y back into the second equation you solved for x: x = 3 + y x = 3 + (0) x = 3 So, y = 0, x = 3.